### v1

# Shear viscosity of quark-gluon plasma at finite temperature and chemical potential and QCD phase transitions

Abstract
We explore the shear viscosity of quark-gluon plasma (QGP) in the full QCD phase diagram within the framework of kinetic theory with the relaxation time approximation based on the $2\leftrightarrow 2$ elastic scatterings of quark quasiparticles. The temperature and chemical potential dependent masses of particles, including $u,d,s$ quarks, their antiparticles, and exchanged mesons, are calculated in the Polyakov-loop extended Nambu–Jona Lasinio (PNJL) model. The results indicate that, at small chemical potential, the value of $\eta/s$ has a minimum near the Mott dissociation of mesons and increases rapidly in the lower-temperature side of the chiral crossover phase transition. At large chemical potential (high density), $\eta/s$ in the QGP phase is dominated by the temperature, and the value of $\eta/s$ is greatly enhanced at lower temperature. At intermediate temperature and chemical potential near the QCD phase transition, the situation is relatively complicated. The behavior of $\eta/s$ is influenced by the competition between temperature, density effect, and QCD phase transition.
I introduction
Understanding the nature of quark-gluon plasma and the phase transition to hadronic matter is presently one of the most concerned questions in quantum chromodynamics (QCD). The calculation from lattice QCD and hadron resonance gas model indicates that the QCD phase transition is a smooth crossover at small chemical potential Aoki et al. (2006); Gupta et al. (2011); Bazavov et al. (2012); Borsányi et al. (2013); Bazavov et al. (2014, 2017); Borsányi et al. (2014); Borsanyi et al. (2020). The simulation of lattice QCD is limited at large baryon chemical potential due to the fermion-sign problem. Many QCD inspired approaches/models predict that a first-order phase transition with a second-order critical end point (CEP) exists at finite temperature and chemical potential, connecting with the crossover transformation (e.g., Fukushima (2008); Ratti et al. (2006); Meisinger and Ogilvie (1996); Costa et al. (2010); Sasaki et al. (2012); Ferreira et al. (2014); Schaefer et al. (2010); Skokov et al. (2011); Qin et al. (2011); Gao et al. (2016); Fischer et al. (2014); Maslov and Blaschke (2023); Fu et al. (2020); Shao et al. (2018); He et al. (2022); Liu et al. (2021)).
The theoretical description of the dynamical evolution of QGP in heavy-ion collisions is needed to build the connection with experimental data. The relativistic viscous hydrodynamics provides an effective framework to describe the collective behaviors and transport properties of QGP Israel and Stewart (1979); Baier et al. (2008); Denicol et al. (2012); Jeon and Heinz (2015). The shear $\eta$ and bulk $\zeta$ viscosities are parameters that quantify the dissipation processes in the evolution of fluid. They reflect the deviation of a medium from an ideal fluid. The simulations indicate that the final hadronic observables are sensitive to the two viscous parameters Romatschke and Romatschke (2007); Song and Heinz (2008); Denicol et al. (2009).
In the earlier studies, the ratios of shear and bulk viscosity to entropy density $\eta/s$ and $\zeta/s$ were taken as constants to simulate the evolution of QGP Heinz and Snellings (2013); Gale et al. (2013). One crucial achievement from the combination of heavy-ion collisions and the viscous hydrodynamics was the discovery that the QGP is a nearly perfect fluid Teaney (2003); Shuryak and Zahed (2004); Shuryak (2005) with the shear viscosity to entropy density ratio $\eta/s$ at very small chemical potential close to $1/{4\pi}$, the KSS limit derived in anti–de Sitter space/conformal field theory (AdS/CFT) correspondence Kovtun et al. (2005); Policastro et al. (2001). Subsequently, it was argued that the value of $\eta/s$ is sensitive to the phase transitions of QGP medium. Similar to water, helium and nitrogen, the shear viscosity to entropy ratio $\eta/s$ of QGP exhibits a minimum at the transition from hadrons to quarks and gluons Csernai et al. (2006); Lacey et al. (2007). Recently, the Bayesian statistical analysis has been adopted to constrain the temperature dependent behavior of $\eta/s$ and $\zeta/s$ at vanishing or small chemical potential Bernhard et al. (2019, 2016); Everett et al. (2021); Nijs et al. (2021); Shen et al. (2023).
The beam energy scan (BES) program at RHIC with the collision energy from 3 to 200 GeV provides a great opportunity to explore the evolution and phase transition of QGP in a wide range of temperature and baryon chemical potential Luo and Xu (2017). One important goal of the BES program is to look for the CEP and the boundary of the first-order chiral phase transition at finite temperature and chemical potential. There are also other planned projects to explore the phase boundary of strongly interacting matter at lower temperature and higher density at the Facility for Antiproton and Ion Research (FAIR) in Darmstadt, the Japan Proton Accelerator Research Complex (J-PARC), the Nuclotron-based Ion Collider Facility (NICA) in Dubna and the High Intensity heavy-ion Accelerator Facility (HIAF) in Huizhou.
The hydrodynamic simulations of the evolution of hot and dense QGP in the corresponding energy regime will play an important role in the analysis of hadronic observables in experiments, and will be helpful to diagnose the QCD phase transition. As the major input parameters, the $T$ and $\mu$ dependent $\eta/s$ and $\zeta/s$ are required in the simulation of hydrodynamical evolution of QGP. Since the direct calculation of these transport coefficients from the non-perturbative QCD on lattice is difficult Kovtun et al. (2005); Astrakhantsev et al. (2017), a growing number of studies have explored them in effective models at vanishing or small chemical potential, such as the hadron resonance gas (HRG) model Noronha-Hostler et al. (2009); Tiwari et al. (2012); Noronha-Hostler et al. (2012); Kadam and Mishra (2015), color string percolation model Sahu and Sahoo (2021), linear sigma model Heffernan et al. (2020), quasiparticle models Mitra and Chandra (2017); Moreau et al. (2019); Soloveva et al. (2020); Mykhaylova et al. (2019); Mykhaylova and Sasaki (2021), Green–Kubo formalism Plumari et al. (2012); Kharzeev and Tuchin (2008); Harutyunyan and Sedrakian (2017); Czajka and Jeon (2017); Gubser et al. (2008); Li et al. (2015), transport approach Wesp et al. (2011); Ozvenchuk et al. (2013); Rose et al. (2018), and QCD motivated model Haas et al. (2014); Christiansen et al. (2015); Dubla et al. (2018); Ghiglieri et al. (2018); Danhoni and Moore (2023); Gao and Liu (2018); Zhang and Hou (2018); Chen et al. (2010, 2013); Gao and Liu (2018).
With the progress of BES program at RHIC, more attentions are paid on the transport coefficients of QGP at finite densities (e.g., Reichert et al. (2021); Shen et al. (2023); Deng et al. (2023) and references therein). To extract the transport coefficients at finite chemical potential where a first-order phase transition and/or the critical region may be involved, a feasible method is taking the calculation in effective quark models describing well both the chiral and confinement phase transitions of QCD. In Ref. Sasaki and Redlich (2010), the shear and bulk viscosities of QGP at finite temperature and chemical potential were derived in the two-flavor quark model. Then the shear viscosity and electric conductivity of hot and dense QGP were calculated in the three flavor quark model Marty et al. (2013); Soloveva et al. (2021). In Ref. McLaughlin et al. (2022), to study the shear viscosity in the presence of a crossover phase transition at finite chemical potential, a combination of the interacting hadron resonance gas in the confined phase and the QCD-based assumption in the deconfined phase was constructed McLaughlin et al. (2022); Borsányi et al. (2018). Note that there are also studies in the NJL-type model with a focus on zero or small chemical potential (e.g., Zhuang et al. (1995); Rehberg et al. (1996a); Xiao et al. (2014)). For more studies related to transport coefficients at finite chemical potential, one can also refer to the researches (e.g., Deb et al. (2016); Soloveva et al. (2020); Kadam et al. (2021)).
An important issue that deserves further exploration is the viscosity of QGP at finite temperature and density as well as its relation with the QCD phase transition. In this work, we will elaborate on the shear viscosity of QGP in the full QCD phase diagram in the framework of three-flavor PNJL quark model. The calculation will be performed within the kinetic theory with the relaxation time approximation based on the $2\leftrightarrow 2$ elastic scatterings between $u,d,s$ quarks and their antiparticles. To derive the cross section and flavor dependent relaxation time, the medium dependent masses of quasiparticles will be calculated. The relationship between $\eta/s$ and the meson Mott dissociation as well as the chiral phase transition, including the critical region and the first-order phase transition with the associated spinodal structure, will be explored in detail.
The paper is organized as follows. In Sec. II, we introduce the 2+1 flavor PNJL quark model and the formulas to calculate the shear viscosity in kinetic theory with the relaxation time approximation. In Sec. III, we illustrate the numerical results of cross section, flavor-dependent relaxation time, the shear viscosity to entropy density ratio, and discuss their relations with the QCD phase transitions. A summary is given in Sec. IV.
II three-flavor PNJL quark model and formula of shear viscosity
II.1 The 2+1 flavor PNJL quark model
We employ the 2+1 flavor PNJL to describe the properties of QGP medium. The Lagrangian density is written as
[…]
$-K\big{[}\texttt{det}_{f}(\bar{\psi}(1+\gamma_{5})\psi)+\texttt{det}_{f}(\bar{\psi}(1-\gamma_{5})\psi)\big{]}$
$-U(\Phi[A],\bar{\Phi}[A],T),$
where $\psi$ denotes the quark fields with three flavors, $u,\ d$, and $s$; $\hat{m}_{0}=\texttt{diag}(m_{u},\ m_{d},\ m_{s})$ in flavor space; $G$ and $K$ are the four-point and six-point interacting constants, respectively. The $\hat{\mu}=diag(\mu_{u},\mu_{d},\mu_{s})$ are the quark chemical potentials. The covariant derivative is defined as $D_{\mu}=\partial_{\mu}-iA_{\mu}$. The gluon background field $A_{\mu}=\delta_{\mu}^{0}A_{0}$ is supposed to be homogeneous and static, with $A_{0}=g\mathcal{A}_{0}^{\alpha}\frac{\lambda^{\alpha}}{2}$, where $\frac{\lambda^{\alpha}}{2}$ is $SU(3)$ color generators.
The effective potential $U(\Phi[A],\bar{\Phi}[A],T)$ is expressed with the traced Polyakov loop $\Phi=(\mathrm{Tr}_{c}L)/N_{c}$ and its conjugate $\bar{\Phi}=(\mathrm{Tr}_{c}L^{\dagger})/N_{c}$. The Polyakov loop $L$ is a matrix in color space
[…]
where $\beta=1/T$ is the inverse of temperature and $A_{4}=iA_{0}$. The Polyakov-loop effective potential Roessner et al. (2007) taken in this study is
$\frac{U(\Phi,\bar{\Phi},T)}{T^{4}}$ $=$ $-\frac{a(T)}{2}\bar{\Phi}\Phi+b(T)\,\mathrm{ln}\big{[}1-6\bar{\Phi}\Phi$
$+4(\bar{\Phi}^{3}+\Phi^{3})-3(\bar{\Phi}\Phi)^{2}\big{]},$
where
$\!a(T)\!=\!a_{0}\!+\!a_{1}\big{(}\frac{T_{0}}{T}\big{)}\!+\!a_{2}\big{(}\frac{T_{0}}{T}\big{)}^{2}\,\,\,\texttt{and}\,\,\,\,\,b(T)\!=\!b_{3}\big{(}\frac{T_{0}}{T}\big{)}^{3},$ (4)
with $a_{0}=3.51$, $a_{1}=-2.47$, $a_{3}=15.2$, $b_{3}=-1.75$ and $T_{0}=210$ MeV. The physical quark mass can be derived in the mean field approximation as
$M_{i}=m_{i}-4G\phi_{i}+2K\phi_{j}\phi_{k}\ \ \ \ \ \ (i\neq j\neq k),$ (5)
where $\phi_{i}$ stands for the quark condensate of the flavor $i$.
The mesons are constructed with the quark-antiquark effective interaction. The pseudoscalar mesons $(\pi^{0},\pi^{\pm},K^{0},\overline{K^{0}},K^{\pm},\eta,\eta^{\prime})$, and the scalar meson partners $\left(\sigma_{\pi^{0}},\sigma_{\pi^{\pm}},\sigma_{K^{0}},\sigma_{\overline{K^{0}}},\sigma_{K^{\pm}},\sigma,\sigma^{\prime}\right)$ are considered to describe the elastic scatterings between quarks and antiquarks. In Eq. (1) the six-fermion interaction in the determinant term can be reduced to an effective four-point interaction in the mean-field approximation, by contracting out $\bar{\psi}\psi$ pairs. For $u$ and $d$ quarks, we currently consider the $SU(2)$ isospin symmetry with the degeneracy condition $m_{u}=m_{d}$ and $\mu_{u}=\mu_{d}$, the effective Lagrangian can be then derived as
$\mathcal{L}=$ $\bar{\psi}(i\gamma^{\mu}D_{\mu}\!+\!\gamma_{0}\hat{\mu}\!-\!\hat{m}_{0})\psi-U(\Phi[A],\bar{\Phi}[A],T)$ (6)
$+\sum_{a=0}^{8}\left[K_{a}^{-}\left(\bar{\psi}\lambda^{a}\psi\right)^{2}+\right.\left.K_{a}^{+}\left(\bar{\psi}i\gamma_{5}\lambda^{a}\psi\right)^{2}\right]$
$+K_{08}^{-}\left[\left(\bar{\psi}\lambda^{8}\psi\right)\left(\bar{\psi}\lambda^{0}\psi\right)+\left(\bar{\psi}\lambda^{0}\psi\right)\left(\bar{\psi}\lambda^{8}\psi\right)\right]$
$+K_{08}^{+}[\left(\bar{\psi}i\gamma_{5}\lambda^{8}\psi\right)\left(\bar{\psi}i\gamma_{5}\lambda^{0}\psi\right)$
$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,+\left(\bar{\psi}i\gamma_{5}\lambda^{0}\psi\right)\left(\bar{\psi}i\gamma_{5}\lambda^{8}\psi\right)].$
The coupling constants in above equation fulfills the following relations
$K_{0}^{\pm}$ $=G\mp\frac{1}{3}K({\cal G}^{u}+{\cal G}^{d}+{\cal G}^{s}),$ (7)
$K_{1}^{\pm}$ $=K_{2}^{\pm}=K_{3}^{\pm}=G\pm\frac{1}{2}K{\cal G}^{s},$
$K_{4}^{\pm}$ $=K_{5}^{\pm}=K_{6}^{\pm}=K_{7}^{\pm}=G\pm\frac{1}{2}K{\cal G}^{u},$
$K_{8}^{\pm}$ $=G\pm\frac{1}{6}K(2{\cal G}^{u}+2{\cal G}^{d}-{\cal G}^{s}),$
$K_{08}^{\pm}$ =±$=\pm\frac{1}{12}\sqrt{2}K({\cal G}^{u}+{\cal G}^{d}-{2\cal G}^{s}),$
where
$\mathcal{G}^{f}=-\frac{N_{c}}{4\pi^{2}}A_{f}(T,\mu),$ (8)
and
$A_{f}(T,\mu)=16\pi^{2}\int_{0}^{\Lambda}\frac{d^{3}p}{(2\pi)^{3}}\left[1-\frac{n_{f}^{+}}{2E}\right.\left.-\frac{n_{f}^{-}}{2E}\right].$ (9)
The $n_{f}^{+}$ and $n_{f}^{-}$ are the modified effective Fermion distribution functions of quark and antiquark with
$\!\!\!n^{+}_{f}\!=\!\frac{\!\Phi e^{\!-\!(\!E_{i}\!-\!\mu_{i}\!)\!/\!T}\!+\!2\bar{\Phi}e^{\!-\!2(\!E_{i}\!-\!\mu_{i}\!)\!/\!T}+e^{-3(E_{i}-\mu_{i})/T}}{\!1\!+\!3\Phi e^{\!-\!(\!E_{i}\!-\!\mu_{i}\!)\!/\!T}\!+\!3\bar{\Phi}e^{\!-\!2(\!E_{i}\!-\!\mu_{i})/T}\!+\!e^{\!-\!3(E_{i}\!-\!\mu_{i})\!/\!T}}$ (10)
and
$\!\!\bar{n}_{f}^{-}\!=\!\frac{\!\bar{\Phi}e^{\!-\!(\!E_{i}\!+\!\mu_{i}\!)\!/\!T}\!+\!2\Phi e^{\!-\!2(\!E_{i}\!+\!\mu_{i}\!)\!/\!T}+e^{-3(E_{i}+\mu_{i})/T}}{\!1\!+\!3\bar{\Phi}e^{\!-\!(\!E_{i}\!+\!\mu_{i}\!)\!/\!T}\!+\!3\Phi e^{\!-\!2(\!E_{i}\!+\!\mu_{i}\!)/T}\!+\!e^{\!-\!3(\!E_{i}\!+\!\mu_{i})\!/\!T}},$ (11)
[…]
With the isospin symmetry of $u$ and $d$ quark, the $\pi^{0}$ is decoupled from the $\eta$ and $\eta^{\prime}$, and the $\pi^{\pm},\pi^{0}$ are degenerate. Similarly, $K_{4}^{\pm}=K_{6}^{\pm}$, which means that the neutral and charged kaons have the same mass. The quark-quark scattering amplitude will be calculated in the random phase approximation, which leads to
$\mathcal{D}_{\pi(\sigma_{\pi})}=\frac{2K_{3}^{\pm}}{1-4K_{3}^{\pm}\Pi_{q\bar{q}}^{P(S)}\left(p_{0},\vec{p}\right)}$ (12)
for the pion ($\sigma_{\pi}$) propagator and
$\mathcal{D}_{K(\sigma_{K})}=\frac{2K_{4}^{\pm}}{1-4K_{4}^{\pm}\Pi_{q\bar{s}}^{P(S)}\left(p_{0},\vec{p}\right)}$ (13)
for the kaon ($\sigma_{K}$) propagator. The label $q$ stands for $u$ and $d$ quark in Eqs. (12)-(13).
In the rest system of reference, the pion mass $M_{\pi}$ and the kaon mass $M_{K}$ satisfy the following relations
$01-4K_{3}^{+}\Pi_{u\bar{u}}^{P}\left(M_{\pi}-i\frac{\Gamma_{\pi}}{2},0\right)=$ 0 (14)
and
$1-4K_{4}^{+}\Pi_{u\bar{s}}^{P}\left(M_{K}-i\frac{\Gamma_{K}}{2},0\right)=0,$ (15)
where $\Gamma_{\pi}$ and $\Gamma_{K}$ are the decay width of pion and kaon, respectively. With the increase of temperature, when the value of $\Gamma_{\pi}$ or $\Gamma_{K}$ changes from zero to nonzero, the dissociation from mesons to quarks occurs, which is also called the Mott phase transition. The polarization function $\Pi_{f\bar{f}^{\prime}}^{P(S)}\left(p_{0},\vec{p}\right)$ is
$\Pi_{f\bar{f}^{\prime}}^{P(S)}\left(p_{0},\mathbf{k}\right)\!=$ $\!\!\!\!-\frac{N_{c}}{8\pi^{2}}\left\{A\left(M_{f},\mu_{f},T\right)+A\left(M_{f^{\prime}},\mu_{f^{\prime}},T\right)\right.$ (16)
$+\left[\left(M_{f}\mp M_{f^{\prime}}\right)^{2}-\left(p_{0}+\mu_{f}-\mu_{f^{\prime}}\right)^{2}+\mathbf{p}^{2}\right]$
×$\left.\times B_{0}\left(|\mathbf{p}|,M_{f},\mu_{f},M_{f^{\prime}},\mu_{f^{\prime}},p_{0},T\right)\right\},$
where $B_{0}\left(|\mathbf{p}|,M_{f},\mu_{f},M_{f^{\prime}},\mu_{f^{\prime}},p_{0},T\right)$ is the source of the imaginary part of the polarization function and the meson mass. It can be written as
$B_{0}\left(|\mathbf{p}|,M_{f},\mu_{f},M_{f^{\prime}},\mu_{f^{\prime}},p_{0},T\right)=$ $\tilde{B}_{0}^{+}\left(-\lambda,|\mathbf{p}|,M_{f},M_{f^{\prime}},\mu_{f},\beta,\varphi\right)-\tilde{B}_{0}^{-}\left(\lambda,|\mathbf{p}|,M_{f},M_{f^{\prime}},\mu_{f},\beta,\varphi\right)$ (17)
$+\tilde{B}_{0}^{+}\left(\lambda,|\mathbf{p}|,M_{f^{\prime}},M_{f},\mu_{f^{\prime}},\beta,\varphi\right)-\tilde{B}_{0}^{-}\left(-\lambda,|\mathbf{p}|,M_{f^{\prime}},M_{f},\mu_{f^{\prime}},\beta,\varphi\right),$
where
$\tilde{B}_{0}^{\pm}\left(\lambda,|\mathbf{p}|,M_{1},M_{2},\mu_{1},\mu_{2},T,\varphi\right)=2\int_{M_{1}}^{\sqrt{M_{1}^{2}+\Lambda^{2}}}\texttt{d}E\,\,k\,n^{\pm}_{f}\,\int_{-1}^{+1}\frac{1}{\lambda^{2}+2\lambda E+2k|\mathbf{p}|x-|\mathbf{p}|^{2}+M_{1}^{2}-M_{2}^{2}}\text{d}x,$ (18)
in which $\lambda=p_{0}+\mu_{1}-\mu_{2}$ and $E=\sqrt{k^{2}+M_{1}^{2}}$.
For the $\eta$ and $\eta^{\prime}$, the quark-antiquark scattering matrix involves the mixing term
$\mathcal{D}_{\eta,\eta^{\prime}}=2\frac{\mathcal{K}}{\mathcal{AC}-\mathcal{B}^{2}}\left(\begin{array}[]{ll}\mathcal{A}&\mathcal{B}\\ \mathcal{B}&\mathcal{C}\end{array}\right)$ (19)
with
$\mathcal{A}=K_{0}^{+}-\frac{4}{3}\mathcal{K}\left(\Pi_{u\bar{u}}^{P}+2\Pi_{s\bar{s}}^{P}\right)$ (20)
$\mathcal{B}=K_{08}^{+}+\frac{4}{3}\sqrt{2}\mathcal{K}\left(\Pi_{u\bar{u}}^{P}-\Pi_{s\bar{s}}^{P}\right)$
$\mathcal{C}=K_{8}^{+}-\frac{4}{3}\mathcal{K}\left(2\Pi_{u\bar{u}}^{P}+\Pi_{s\bar{s}}^{P}\right)$
$\mathcal{K}=K_{0}^{+}K_{8}^{+}-K_{08}^{2}$
In the numerical calculation, a cut-off $\Lambda$ is implemented in the three-momentum space for divergent integrations. We take the model parameters obtained in Rehberg et al. (1996b): $\Lambda=602.3$ MeV, $G\Lambda^{2}=1.835$, $K\Lambda^{5}=12.36$, $m_{u,d}=5.5$ and $m_{s}=140.7$ MeV, determined by fitting $f_{\pi}=92.4$ MeV, $M_{\pi}=135.0$ MeV, $m_{K}=497.7$ MeV and $m_{\eta}=957.8$ MeV.
II.2 Shear viscosity in kinetic theroy with the relaxation time approximation
We shall calculate the transport coefficients in the PNJL model based on the kinetic theory with the relaxation time approximation. The quarks, antiquarks and mesons are taken as quasiparticle in the elastic scatterings. It has been found that the transport coefficients derived in kinetic theory with the quasiparticle approximation are close to the one-loop diagram calculations based on Kubo relations. In this study, we will focus on the properties of shear viscosity at different temperature and baryon density.
The shear viscosity from the kinetic theory with relaxation time approximation Hosoya and Kajantie (1985); Gavin (1985); Sasaki and Redlich (2009, 2010); Chakraborty and Kapusta (2011); Albright and Kapusta (2016) are written as
$\eta=\frac{1}{15T}\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\frac{p^{4}}{E_{\alpha}^{2}}d_{q}\left(\tau_{\alpha}f_{\alpha}^{0}(1-f_{\alpha}^{0})\right)$ (21)
where the subscript $\alpha$ stands for quarks ($u,d,s$) and antiquarks ($\bar{u},\bar{d},\bar{s}$) and the degeneracy factor $d_{q}=2N_{c}$ for spin and color. The relaxation time $\tau_{\alpha}$ for particles $\alpha$ depends on the environment of QGP medium, i.e., it varies with temperature and chemical potential. $f^{0}_{\alpha}\left(E_{\alpha},T,\mu_{\alpha}\right)$ describes the distribution function in local equilibrium for quarks and antiquarks. The effective distribution function given in Eqs. (10) and  (11) will be used in the specific calculation.
Eq. (21) indicates that the relaxation time plays a very important role in calculating the shear viscosity coefficient. It is related to the microscopic scattering process between particles in the thermal medium. In this work, the two-body elastic scatterings, including the quark-quark, quark-antiquark and antiquark-antiquark, are considered to calculate the shear viscosity of QGP. The averaged relaxation time of particle species $i$ can be derived as
$\tau_{i}^{-1}\left(T,\mu_{q}\right)=\sum_{j=q,\bar{q}}\rho_{j}\left(T,\mu_{q}\right)\bar{w}_{ij},$ (22)
where $\rho_{j}\left(T,\mu_{q}\right)$ is the quark or antiquark number density of species $j$. $\bar{w}_{ij}$ is the averaged transition rate defined as
$\bar{w}_{ij}=$ $\frac{1}{\rho_{i}\rho_{j}}\int\frac{d^{3}\mathbf{p}_{i}}{(2\pi)^{3}}\int\frac{d^{3}\mathbf{p}_{j}}{(2\pi)^{3}}d_{q}f_{i}^{(0)}\left(E_{i},T,\mu_{q}\right)d_{q}$ (23)
×$\times f_{j}^{(0)}\left(E_{j},T,\mu_{q}\right)\cdot v_{\mathrm{rel}}\cdot\sigma_{ij\rightarrow cd}\left(s,T,\mu_{q}\right).$
Here $v_{\mathrm{rel}}$ is the relative velocity and $\sqrt{s}$ is the center-of-mass energy. $\sigma_{ij\rightarrow cd}$ is the cross section from the initial incident (anti)quarks $(i,j)$ to the final outgoing (anti)quarks $(c,d)$ at collision energy $\sqrt{s}$.
The relative velocity $v_{\mathrm{rel}}$ can be calculated in the c.m frame with
$v_{\text{rel }}=\frac{\sqrt{\left(E_{i}^{*}E_{j}^{*}-\left|\mathbf{p}_{i}^{*}\right|\left|\mathbf{p}_{j}^{*}\right|\right)^{2}-\left(M_{i}M_{j}\right)^{2}}}{E_{i}^{*}E_{j}^{*}}$ (24)
where
$E_{i}^{*}=\frac{s+M_{i}^{2}-M_{j}^{2}}{2\sqrt{s}},\,\,\,\,E_{j}^{*}=\frac{s-M_{i}^{2}+M_{j}^{2}}{2\sqrt{s}}$ (25)
and
$\left|\mathbf{p}_{i}^{*}\right|=\frac{\sqrt{\left(s-\left(M_{i}+M_{j}\right)^{2}\right)\cdot\left(s-\left(M_{i}-M_{j}\right)^{2}\right)}}{2\sqrt{s}}.$ (26)
The total cross section can be derived with the integral of the differential cross section
$\sigma=\int_{t_{-}}^{t^{+}}\mathrm{d}t\frac{\mathrm{d}\sigma}{\mathrm{d}t}\left(1-f^{(0)}(E_{c}^{*},T,\mu_{q})\right)\left(1-f^{(0)}(E_{d}^{*},T,\mu_{q})\right)$ (27)
where the factor $(1-f^{(0)})$ represents the Pauli blocking for quarks and/or antiquarks in the final state; $t$ is the Mandelstam variable; $t^{\pm}$ indicates the upper and lower limits of $t$
$t_{\pm}=$ $M_{i}^{2}+M_{c}^{2}-\frac{1}{2s}\left(s+M_{i}^{2}-M_{j}^{2}\right)\left(s+M_{c}^{2}-M_{d}^{2}\right)$ (28)
±$\!\pm\!2\!\left.\sqrt{\!\frac{\left(s\!+\!M_{i}^{2}\!-\!M_{j}^{2}\right)^{2}}{4s}\!-\!M_{i}^{2}}\right.\!\sqrt{\frac{\left(s\!+\!M_{c}^{2}\!-\!M_{d}^{2}\right)^{2}}{4s}\!-\!M_{c}^{2}}.$
The differential cross section in Eq. (27) is written as
$\frac{d\sigma}{dt}=\frac{1}{16\pi s_{ij}^{+}s_{ij}^{-}}\frac{1}{4N_{c}^{2}}\sum_{sc}|\mathcal{M}|^{2},$ (29)
where the $s_{ij}^{\pm}$ is defined as
$s_{ij}^{\pm}=s-\left(M_{i}\pm M_{j}\right)^{2}.$ (30)
$\frac{1}{4N_{c}^{2}}|{\mathcal{M}}|^{2}$ denotes the matrix element squared averaged over the color and spin of the incident particles, and summed over the final scattered particles. The calculation of $|{\mathcal{M}}|^{2}$ is related to the specific scattering channels, which can be performed using the standard methods in quantum field theory. One can refer to Refs. Rehberg et al. (1996a); Soloveva et al. (2021) for more details.
For the quark-quark scattering, only $u$ and $t$ channel are involved. The mesons exchanged in different scattering processes are listed in Table. 1. On the premise of isospin symmetry of $u,d$ quark, the process of $u+s\rightarrow u+s$ is equivalent to that of $d+s\rightarrow d+s$, and the process of $u+u\rightarrow u+u$ is equivalent to that of $d+d\rightarrow d+d$. For the quark-antiquark scattering, only $t$ and $s$ channel are involved. The mesons exchanged in different processes are listed in Table. 2. The antiquark-antiquark scattering can be calculated using the same steps as the quark-quark scattering.
Table 1: Mesons exchanged in the $t$ and $u$ channels in the different quark-quark scattering processes.
process
exchanged mesons
($u$ channel)
exchanged mesons
($t$ channel)
$ud\rightarrow ud$ $\pi,\sigma_{\pi}$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$
$uu\rightarrow uu$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$
$us\rightarrow us$ $K,\sigma_{K}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$
$ss\rightarrow ss$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$
Table 2: Mesons exchanged in the $s$ and $t$ channels for in the different quark-antiquark scattering processes.
process
exchanged mesons
($s$ channel)
exchanged mesons
( $t$ channel)
$u\bar{d}\rightarrow u\bar{d}$ $\pi,\sigma_{\pi}$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$
$u\bar{u}\rightarrow u\bar{u}$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$
$u\bar{u}\rightarrow d\bar{d}$ $\pi,\eta,\eta^{\prime},\sigma_{\pi},\sigma,\sigma^{\prime}$ $\pi,\sigma_{\pi}$
$u\bar{s}\rightarrow u\bar{s}$ $K,\sigma_{K}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$
$u\bar{u}\rightarrow s\bar{s}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$ $K,\sigma_{K}$
$s\bar{s}\rightarrow u\bar{u}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$ $K,\sigma_{K}$
$s\bar{s}\rightarrow s\bar{s}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$ $\eta,\eta^{\prime},\sigma,\sigma^{\prime}$
III Numerical results and discussion
In this section, we present the numerical results, including the Mott dissociation of mesons, the cross section, the flavor dependent relaxation time and the shear viscosity to entropy density ratio in the full QCD diagram. In particular, we will discuss the dependence of the relaxation processes and the transport coefficient on temperature and chemical potential as well as the QCD phase transition.
III.1 Mott dissociation of mesons
The dissociation of a meson involves its physical mass and constitute mass in the QGP medium. When the physical mass of a bound meson equals the constituent one, the meson becomes unstable and undergoes a Mott dissociation (Mott phase transition) Blaschke et al. (2017). The corresponding temperature is referred to as the Mott phase transition temperature, $T_{\text{Mott}}$. Above the $T_{\text{Mott}}$, mesons will dissociate into quark-antiquark pairs. From the onset of Mott phase transition the meson decay width becomes non-zero and increases with rising temperature.
Figure 1: Masses of pseudoscalar mesons (pion and kaon), the constitute masses $M_{u}+M_{d}$ and $M_{u}+M_{s}$, and the meson decay widths as functions of temperature for $\mu_{B}=0$ (upper) and $\mu_{B}=600$ MeV (lower).
Figure 1 presents the physical masses, constitute ones and decay widths of pseudoscalar mesons (pion and kaon) as functions of temperature at $\mu_{B}=0$ and $\mu_{B}=600$ MeV. It shows that the constitute mass $M_{u}+M_{d}$ ($M_{u}+M_{s}$) of pion (kaon) decreases with increasing temperature due to the partial restoration of chiral symmetry. However, the physical mass of bound pion (kaon) derived from Eqs. (14) and (15) increases with rising temperature. Fig. 1 also indicates that the Mott phase transition at $\mu_{B}=600$ MeV is lower than that at $\mu_{B}=0$. This is because that not only the temperature but also the density effect can cause the restoration of chrial symmetry. It can be also observed that the pion and kaon undergo the Mott dissociations at nearly the same but slightly different temperature.
Figure 2: Mott phase transitions of pion and kaon and chiral phase transition of light $u,\,d$ quarks in the $T-\mu_{B}$ plane. The isentropic trajectories for $s/\rho_{B}=300,100,50,25,15,10,5,4,3$ are also plotted for the convenience of the subsequent discussion.
Figure 3: Mott phase transitions of pion and kaon and chiral phase transition of light $u,\,d$ quark in the $T-\rho_{B}$ plane.
To indicate the relationship between the Mott dissociation of pseudoscalar mesons and the chiral phase transition of quarks, we draw in Figs. 2 and 3 these phase transition lines in the full QCD phase diagram. Fig. 2 demonstrates that for each given chemical potential the temperature of Mott phase transition is close to but slightly higher than that of chiral crossover transition of $u$ ($d$) quark. The difference between them is less than $10$ MeV in the whole crossover region, and continually decrease when approaching the CEP. Although these phase transition lines are close to each other in the $T-\mu_{B}$ plane near the critical region, there is a larger difference in the temperature-density plane, as demonstrated in Fig. 3. This attributes to that the baryon density in the critical region being sensitive to the chemical potential. Fig. 3 also indicates that the Mott dissociation occurs in the metastable phase for pion and both the metastable and unstable phases for kaon in the first-order phase transition range at lower temperature. The isentropic trajectories are complicated in the spinodal region as shown in Fig. 2, which implies that the thermodynamic and transport properties of QGP are complex in the range of the first-order phase transition when the spinodal structure is considered.
III.2 Shear viscosity
We present in Fig. 4 the cross sections of quark-quark and quark-antiquark scatterings at different temperature as functions of $\sqrt{s}$ for $\mu_{B}=0$ and $\mu_{B}=800$ MeV. For each scattering process, it is found that the cross section is relatively larger at the temperatures near the crossover phase transition and Mott phase transition. When the temperature is much higher than that of Mott phase transition, the cross sections decrease since the $\pi$ and $K$ mesons have larger decay widths, as shown in Fig. 4. Fig. 4 also shows that the thresholds of particles in each process change with temperature and chemical potential. Compared with $ud\rightarrow ud$ ($u\bar{d}\rightarrow u\bar{d}$), the threshold of $us\rightarrow us$ ( $u\bar{s}\rightarrow u\bar{s}$) at the same temperature and chemical potential is postponed to a larger $\sqrt{s}$. The reason is that $\sqrt{s}$ should at least equal to the larger of the sum of incident particle mass and the sum of outgoing particle mass, and $M_{u}+M_{s}$ is always larger than $M_{u}+M_{d}$ in the whole QCD phase diagram. Fig. 4 also indicates that the cross sections of quark-antiquark scattering are much larger than those of quark-quark scattering. This mainly attributes to the resonance of exchanged meson with the incident quarks in the $s$ channel in quark-antiquark scattering, which leads to a distinct peak-like structure existing in Fig. 4(c) and (d).
Figure 4: Cross sections of quark-quark (including $ud\rightarrow ud$ and $us\rightarrow us$ ) and quark-antiquark (including $u\bar{d}\rightarrow u\bar{d}$ and $u\bar{s}\rightarrow u\bar{s}$) elastic scattering as functions of center-of-mass collision energy at different temperature and quark chemical potential.
Figure  5 present the relaxation time of different particle species as functions of temperature at $\mu_{B}=0$ and $\mu_{B}=800$ MeV. For the $u,d$ isospin symmetric quark matter, the relaxation time $\tau_{u}=\tau_{d}$ and $\tau_{\bar{u}}=\tau_{\bar{d}}$ always hold. In particular, $\tau_{u}=\tau_{d}=\tau_{\bar{u}}=\tau_{\bar{d}}$ at vanishing chemical potential. In the case of $\mu_{B}=0$ , it can be seen that the relaxation time $\tau_{s}$ of strange quark is larger than that of $u$ ($d$) quark near the region of chiral and mott phase transitions. The chiral crossover transition temperature of $s$ quark is about $250$ MeV at $\mu_{B}=0$. We can see that $\tau_{s}$ gradually approaches $\tau_{u}$ at high temperature after the chiral restoration of $s$ quark, which is just the expected behavior under the SU(3) symmetry. At finite chemical potential, $\tau_{u}\neq\tau_{\bar{u}}$ and $\tau_{s}\neq\tau_{\bar{s}}$ due to the different distribution function of quark and antiquark. Numerically, the relaxation time of quarks are larger than those of antiquarks. Fig. 5 also indicates that the relaxation for each particle species has a concave shape, which is to certain degree connected with the shear viscosity of QGP medium.
Figure 5: Relaxation time of quarks and antiquarks as functions of temperature for $\mu_{B}=0$ and $\mu_{B}=800$ MeV. The solid (hollow) triangles mark the temperatures of chiral (Mott) phase transitions for different particle species.
In the following, we elaborate on the shear viscosity of QGP under different conditions. Fig. 6 presents the shear viscosity to entropy density ratio $\eta/s$ as functions of temperature at different chemical potentials. For $\mu_{B}$=0, the value of $\eta/s$ are very small above the temperature of meson Mott phase transition. The minimum of $\eta/s$ is about 0.1, close to the KSS limit, which indicates that the QGP medium is a nearly perfect” fluid in the extreme relativistic high-energy collision. The value of $\eta/s$ increases with the rising chemical potential. It can be observed from Fig. 6 that $\eta/s$ is slightly larger than 1 near the CEP, and the value will continually increase at lower temperature for larger chemical potentials.
Figure 6: Shear visicosity to entropy ratio $\eta/s$ as functions of temperature at different $\mu_{B}$. The solid (hollow) triangles mark the temperatures of chiral (Mott) phase transitions.
The minimum of $\eta/s$ at $\mu_{B}=0$ is closer to the Mott phase transition (meson dissociation) than the chiral phase transition of $u,d$ quark, and $\eta/s$ increases rapidly with the occurrence of hadronization. In Fig. 6 some values of $\eta/s$ below the chiral phase line are also preserved to display the variation of $\eta/s$ near the chiral crossover phase transition region. However we should keep in mind that the collisions between hadrons dominate the transport process in the hadronic phase at low temperature. It can be observed from Fig. 6 that the location of the minimum $\eta/s$ at finite chemical potential become more and more flat, and the relationship with phase transitions is not as distinct as that of $\mu_{B}=0$.
Fig. 7 describes the behaviors of $\eta/s$ as functions of chemical potential at several temperatures lower than or equal to 250 MeV. For the temperature higher than that of the chiral phase transition of u(d) quark, e.g., $T=220$ or $250$ MeV, it can be seen that $\eta/s$ monotonically increases with rising chemical potential. The nonmonotonic behavior appears with the decrease of temperature, and the location of the minimum of $\eta/s$ gradually approaches to the chiral phase transition (solid triangle in Fig. 7), particularly in the critical region. As indicated in the formula of shear viscosity in Eq. (21), the contribution of density effect becomes more and more important with the rising chemical potential and the decreasing temperature.
Figure 7: $\eta/s$ as functions of baryon chemical potential at different temperatures. The solid triangles (dots) mark the chemical potentials of chiral phase transition in the crossover region (in the first-order region).
Fig. 7 also shows the features of $\eta/s$ at $T=100,75,50$ MeV with inclusion of the first-order phase transition associated with the spinodal structure. The solid dots on each curve with the same color indicates the value of $\eta/s$ on the boundaries of the first-order phase transition. From higher chemical potentials to lower ones, the first solid dot encountered on each curve corresponds to the high-density side of the first-order phase transition, and the second solid dot corresponds to the low-density side. This figure indicates that the values of $\eta/s$ are quite different on the boundaries of the chiral broken and chiral restored phases at lower temperature.
Figure 8: $\eta/s$ as functions of baryon density at different temperatures. The solid triangles (dots) mark the baryon densities of chiral phase transition in the crossover region (in the first-order transition region).
Fig. 8 presents the value of $\eta/s$ in the $T-\rho_{B}$ plane for different temperatures. The dashed curve is the boundary of the first-order phase transition. The behaviors of $\eta/s$ at $T=200$ MeV or higher temperatures are similar to those in the $T-\mu_{B}$ plane plotted in Fig. 7. However, compared to Fig. 7, in the $T-\rho_{B}$ plane the distance between the density of the minimum $\eta/s$ and that of the chiral phase transition appears to be enlarged when approaching to the critical region of the CEP from the high-temperature side. The reason is that the baryon density is sensitive to chemical potential in the critical region. Small variation of chemical potential will lead to a larger change of baryon number density. For the first-order phase transition, the difference of $\eta/s$ on the two phases is small near the CEP. However, it can be seen that at lower temperature, e.g., $T=50$ MeV, $\eta/s$ in the chiral restored QGP phase is much larger than the chiral breaking phase. This figure clearly indicates the shear viscosity to entropy density ratio at high baryon densities highly depends on the temperature. The value of $\eta/s$ increases with the decline of temperature, which is qualitatively consistent with the variation of shear viscosity of common liquids.
Figure 9: The value of $\eta/s$ on the phase boundaries including the chiral corssover and first-order phase transitions.
We plot in Fig. 9 the value of $\eta/s$ on the boundaries of the chiral phase transition of $u,d$ quarks. Fig. 9(a), (b) and (c) present the behavior of $\eta/s$ as functions of temperature, baryon chemical potential and baryon density, respectively. Since the relation between $\rho_{B}$ and $\mu_{B}$ on the crossover phase transition line near the critical region is nonmonotonic as indicated in Fig. 2 and Fig. 3, Fig. 9(c) shows a similar structure for $\eta/s$. For the first-order phase transition, Fig. 9(a) also indicates that $\eta/s$ in the chiral restored QGP phase is smaller than that in the chiral breaking phase in the region not far away from the critical region. This situation reverses as the temperature decreases (correspondingly, the chemical potential increases as shown in Fig. 9(b)). Such a feature reflects that the value of $\eta/s$ at intermediate chemical potential and temperature are affected by the competition between the temperature, chemical potential (density) and QCD phase transition.
Figure 10: $\eta/s$ along the different isentropic trajectories. The isentropic trajectories are plotted in Fig. 1. The solid triangles (dots) mark the temperature of chiral phase transition in the crossover region ( of the higher-density side in the first-order region).
Finally, we demonstrate in Fig. 10 the behavior of $\eta/s$ along the different isentropic trajectories plotted in Fig. 1. It can be seen that $\eta/s$ has a minimum on each isentropic trajectory. The value of $\eta/s$ changes slowly for larger $s/\rho_{B}$ except for the region near the chiral phase transition. This situation corresponds to the fireball generated in high energy collision. For the case of smaller $s/\rho_{B}$, the rapid change of $\eta/s$ occurs near the first-order phase transition with the associated spinodal structure. The numerical results demonstrate that the minimum value of $\eta/s$ on each curve appear near the location of the first-order phase transition, as indicated with the solid dots in Fig. 10. The solid dots mark the boundary of the first-order phase transition at the high-density side. It is helpful to understood the result in conjunction with QCD phase diagrams shown in Fig. 2.
IV summary
In this work, we explored the shear viscosity of QGP in the full QCD phase diagram within the framework of kinetic theory with the relaxation time approximation. The flavor dependent relaxation time is derived according to the $2\leftrightarrow 2$ elastic scattering between quasiparticles including the $u,d,s$ quarks and their antiparticles. The temperature and chemical potential dependent masses of quasiparticles are calculated in the PNJL model. The masses of exchanged mesons are derived with the random phase approximation.
The numerical results show that, for small chemical potential, the relaxation time for each kind of quark and its antiparticle has a minimum near the Mott dissociations of mesons in the high-temperature side. The relaxation time increases rapidly below the chiral crossover phase transition line. Correspondingly, from high temperature to lower one the value of $\eta/s$ has a concave behavior and increase rapidly in the lower-temperature side of the crossover phase transition. The minimum value of $\eta/s$ at $\mu_{B}=0$ is about 0.1, close to the KSS limit and the value extracted from the experimental data.
At large chemical potential (high density), the behavior of $\eta/s$ in the QGP phase is mainly dominated by temperature. The value of $\eta/s$ of dense QGP is greatly enhanced with the decline of temperature. $\eta/s$ is about 1 when approaching to the critical region from the high-temperature side. The situation becomes relatively complicated for the first-order phase transition. The value of $\eta/s$ is influenced by the competition between temperature, density effect and QCD phase transition. The evolving behaviors of $\eta/s$ in the QCD phase diagram can be referenced by hydrodynamics to simulate the QCD phase transition.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grant No. 11875213 and Natural Science Basic Research Plan in Shaanxi Province of China (Program No. 2024JC-YBMS-018).